Ticket Number: HEA-NOC/20040815-1 Ticket Status: SUPERCEDE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20040815 09:59 UTC+1 Problem Start: 20040815 06:56 UTC+1
Ticket Closed: 20040830 10:57 UTC+1 Problem End: 20040815 06:57 UTC+1
Superceded by: 20040519-1
Between 06:56 and 06:57 there were two quick up / downs to UCC's 100M link
20040815 09:56 SOB Issue seems to have resolved its self. No response from
cork. Will contact again on Monday to check for RFO.
20040816 09:31 SOB Line stable since these outages. Closing ticket
20040816 15:59 SOB UCC report that outage was caused by a user scheduling
highly parallelised download at that time, causing an OSPF outage.
Requested confirmation that user has been identifing and stopped.
20040818 14:05 SOB outage re-occured yesterday morning at 0700 and 0800.
User download was not cause this time. Henry is investigating further.
20040823 09:28 JC Taking ownership of the ticket.
20040824 09:21 JC Henry emailed me saying that UCC had eliminated any
Computer Centre controlled scheduled events for those times.
I sent logs for the ATM1/0.4079 interface on 20040817 showing the line
20040824 09:39 JC After suggesting high cpu utilisation as a possible
cause, I recieved a mail from Henry stating that their logs showed no cpu
saturation around the time of the transitions, and their router had the
same down – neighbor down – up sequence in the logs.
20040830 10:56 AH problems tracked in more detail in ticket 20040519-1
This ticket can be monitored at http://www.hea.net/tickets/20040815-1